OpenStudy (anonymous):

what is the sum of the arithmetic sequence 3,9,15, if there are 36 terms

OpenStudy (stacey):

You should have a formula for finding this sum. n is the number of terms, so for this problem n=36. d is the common difference. It can be found by subtracting the second term minus the first.

OpenStudy (anonymous):

i no its 6

OpenStudy (anonymous):

u didn't try the formula i posted in ur previous post?

OpenStudy (anonymous):

i guess not

OpenStudy (stacey):

Then the formula to use is \[(n/2) * [2a_{1} + (n-1)*d]\]

OpenStudy (anonymous):

y?just use the formula n u will get the sum, it needs simple airthmetics using a calculator

OpenStudy (anonymous):

yes stacey is right

OpenStudy (stacey):

\[a_{1} =3\] since 3 is the first term.

OpenStudy (anonymous):

i got 151

OpenStudy (anonymous):

sorry wrong post

OpenStudy (anonymous):

i believe the answer is 4234

OpenStudy (stacey):

That seems too big.

OpenStudy (anonymous):

3468

OpenStudy (anonymous):

3888

OpenStudy (stacey):

n/2 = 36/2 = 18 so we have 18*[2*3 + (36-1) * 6]

OpenStudy (anonymous):

are either of those answers right

OpenStudy (stacey):

3888 is correct.

OpenStudy (anonymous):

3468 or 3888